114583:

If g = 77 cm and h = 85 cm, what is the length of f?

vietress:

Need more context

vietress:

Are you trying to find the length of the hypotenuse?

114583:

a leg.

vietress:

A leg...

vietress:

The height?

114583:

yea the Legnth

vietress:

The formula to find the length of the hypotenuse is \[a^2 + b^2 = c^2\] If you're trying to find one of the other sides (a leg) then instead of plugging in numbers for a and b, you plug in numbers for a OR b, and c. In this case, it would be: \[a^2 + 77^2 = 85^2\] The next step is to solve the exponents. 77 squared is 5929 85 squared is 7225 Now the new equation is: \[a^2 + 5929 = 7225\] Now, you subtract 5929 from both sides to isolate the variable. 7225 - 5929 = 1296 This gives us: \[a^2 = 1296\] The final step is to take the square root of both sides of the equation. \[\sqrt{a^2} = \sqrt{1296}\] The final answer is a (in other words "f") = 36 Hope this helped!

114583:

omg thank youu

vietress:

np :)

vietress:

would you mind marking my answer as "best response"?